E-time Observing - 20170205 UT

Observers: JHill, OKuhn (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: GTaylor, DMiller, XZhang, PHinz
Telescope Support: (LBTO)
Software Support: MDelapena, CBiddick, SHooper (Tucson)
Instrument Support: (LBTO)
Engineering Support:

Summary

TCS tests were successfully completed, although there are still some minor things with GCS to work out later.

AO tests from AOeng account were successfully completed, also with minor things to work out later (repoint fails [differently from LBTI], stopAO from IIF doesn't always open the loop. See the Adaptive Optics wiki log for more details from Doug/Greg http://wiki.lbto.org/bin/view/AdaptiveOptics/20170205_DX . We took some fast tip-tilt data with the AO loop closed with IRTC@RFBG.

Science observations were completed with MODS/MODS MOS for a Potsdam DDT program.

AGw1 probe centering was very strange as we could never get a circle more than ~2 arcsec in diameter. But, the probe seems to be in the correct position with a very small adjustment.

Pupil centering went well after Doug changed IDL code to match the latest GCSL output.

Collecting transform data failed completely due to IDL-Telescope-LUCI interaction issues which may have included inexperienced users (although I claim I've done this before). And was saved by thick clouds which wiped out the data in any case...

Tried NS guiding on moderate rate asteroid at the very end, but there was too much twilight and cloud to find the guidestar.

Preparations

Running AO as AOEng@ for local SVN code in CentOS6.

New AOS version 16.4 with improved error messages.

New GCS version 5.4f. New PCS version, and new PCSGUI.

New mcspu for better log file management

Details

TCS and GCS Testing with overlapping AO Tests

Authorized as LUCI/LUCI (with IRTC2 mounted at RFBG)

02:04 ACTIVE preset for collimation (some coma on DX, but still 100 arcsec from RX limit) Seeing is 0.7 to 1.0 arcsec on guider.

Guiding Hotspot: SX 242,256 DX 275,243

Changing hotspot in Y reported probe at limit (was an on-axis preset)

Testing manual changing of the hotspot on SX. Doesn't seem to display a pause wfs message on GUI, but also doesn't get wfsing stuck. We've not understood the behaviour in real time, but plenty of log messages to analyze.

02:18 ACE_ACTIVE preset

Another hotspot test nn DX with RunAO active....... (guiding and wfs were not sending to PCS and PSF)
Sun Feb 5 02:37:12.674 2017 setHS: 250.0, 243.0 [pix]. (pauses wfs, and then moves probe)
Sun Feb 5 02:38:48.102 2017 setHS: 275.0, 243.0 [pix]. (back to original, but doesn't resume wfs)

After a while, it did resume wfsing........
Michele now has some more log data to analyze the precise behaviour.

SX Sun Feb 5 02:53:41.955 2017 setHS: 275.0, 256.0 [pix].

We adopted the various new TCS patches.

AO Testing

03:21 ACE_ACTIVE preset, RunAO, etc.

04:00 Close side vent doors (reduced TT significantly, but still a lot of TT at 2.5 Hz)

04:18-ish colllecting IRTC2 data at RFBG with AO loop closed. See AO wiki log for filenames and AO data sets. The 2.5 Hz peak was mostly gone during this data. SX OVMS+ TT was 19 mas rms, and DX OVMS+ TT was 25 mas rms.

04:35 DX open-loop data on IRTC2

-rw-r--r--. 1 ups 500   178560 Feb  5 04:09 irtc2.20170205.040909.fits
-rw-r--r--. 1 ups 500 16387200 Feb  5 04:09 irtc2.20170205.00002_cube.fits

-rw-r--r--. 1 ups 500    20160 Feb  5 04:18 irtc2.20170205.041806.fits (apertnr=3 H-band 10000 microsec)
-rw-r--r--. 1 ups 500  4803840 Feb  5 04:19 irtc2.20170205.00003_cube.fits  (48x50, 1000 frames, closed loop)
-rw-r--r--. 1 ups 500    20160 Feb  5 04:20 irtc2.20170205.042002.fits  (apertnr=3 H-band 10000 microsec)
-rw-r--r--. 1 ups 500 48003840 Feb  5 04:34 irtc2.20170205.00004_cube.fits  (48x50, 10000 frames, closed loop)

-rw-r--r--. 1 ups 500    20160 Feb  5 04:35 irtc2.20170205.043513.fits  (apertnr=3 H-band 10000 microsec)
-rw-r--r--. 1 ups 500  4803840 Feb  5 04:36 irtc2.20170205.00005_cube.fits (48x50, 1000 frames, open loop)
-rw-r--r--. 1 ups 500    20160 Feb  5 04:40 irtc2.20170205.044005.fits  (apertnr=3 H-band 10000 microsec)
-rw-r--r--. 1 ups 500  4803840 Feb  5 04:41 irtc2.20170205.00006_cube.fits (48x50, 1000 frames, dark)

04:50 Testing RunAO repoint from AOS command gui for SX. Failed (differently than LBTI).

Testing a LUCI preset

04:56 Set Pseudomono on SX

04:57 ACTIVE preset from Michelle's OT-generated script, but we were expecting ACE_ACTIVE. Then LUCI1 MOS switches to unknown (but turned out to be misinterpretation of GUI).

StopAO from IIF did not work to open the AO loop on the right side.

ACQUIRE preset

05:10 ACTIVE preset again from LUCI - pseudomono back on, so presetting on both sides.

MODS Binocular Science Observations

05:15 Reconfigure to MODS/MODS for science observation.

Problem with swing arm or OSS took an extra 7 minutes.

05:43 ACQUIRE preset with MODS/MODS

05:48 ACTIVE preset with MODS/MODS - seeing 0.8 arcsec on guider at the start. Some coma on DX made collimation take longer. Clear sky for the first long MODS exposure, but thin clouds coming in during the second exposure.

07:58 Begin reconfiguration for LUCI/LUCI.

AGW1 Alignment Activities

AGw1 guide probe centering

08:16 ACQUIRE preset with LUCI/LUCI SX IE=-37 CA=-45 DX IE=+76 CA=-207

08:21 ACTIVE preset BS9148

stop Guiding
rotator to 0 deg
ptincrement 0 CA -5
gcsclient readGuideCam left -e 1000 (first one failed as expected)
gcsclient readGuideCam left -e 90000

data in /lbt/data/share/tcs/telescope.20170205/left_guider000001.fits

08:29 guide 0001
118,106 measured (tiny circle saturated)
236,212 x2 for binning
251,251 corrected for DETSEC
compare to hotspot 242,256
need to decrease ROT homeoffset by 0.45 arcsec = .27 mm = -18 counts 35578 --> 35560
need to decrease RAD oofset by 0.20 arcsec = 0.12 mm = -120 microns 417825 --> 417705

ssh oac@oac
vi oacontrol.conf
rdwrconfig -u 1
stop and restart GCSL

repeat process again , but X was -0.1 mm, Y was 612.50

08:54 guide 0002
120,106 (tiny ellipse saturated)
WHYYYYYYYYYYYYYYYYYYYYYYYYY? is it not a bigger circle

ptincrement 0 CA -10
08:57 guide 0003

Reasons for WHY?
  • moving the left rotator cancelled the left preset (so it did move the LFBG rotator).
  • Geno and I both watched the LFBG rotator moving on MCSGUI.
  • ptincrement 0 CA 2 did perturb the guide loop when the loop was closed.

guide 0004 is a bit of a circle, but with collimation problems.

Didn't figure out the problem, but moving on to the next step anyway.

AGW1 Off-axis Pupil Wobble measurements

09:15 pupil_collect, retval, EXP=10.0
   Acquiring LEFT WFSC image (exp=10.0000 s)...
% Attempt to subscript RESULT with  is out of range.
% Execution halted at: GCS_TAKE_IMAGE    169
   /home/lbto/idl/wfsc/gcs/gcs_take_image.pro
%                      COLLIMATE_ONCE     50
   /home/lbto/idl/wfsc/collimate/collimate_once.pro
%                      PUPIL_COLLECT      70
   /home/lbto/idl/wfsc/pupil/pupil_collect.pro
%                      $MAIN$          

09:18 repeated pupil_collect with same message again.

Called DougM -

First problem seems to be a different output format when asking gcs to take an image from IDL. idl/wfsc/gcs/gcs_take_image.pro
(Did this change since September 2016?)

Second minor confusion seems to be that working files are in ../telescope.20170205/ (the working directory where I ran IDL) while images are in ../lbto.20170205/.

Now using /lbt/data/share/tcs/lbto.20170205

09:45 clouds overhead getting heavier

09:53 pupil_collect now running (by Doug as LBTO) (round 1)

Data files temporarily in /home/lbto/ moved later to /home/lbto/data/20170205_lfbg/

Block sending centroids to PCS
Move M3 Tip/Tilt in Global Offset Increments of 50 arsec
Recover pointing with IE/CA from pngui Paddle
Resume sending centroids to PCS
stop WFSing
Do another round of pupil_collect.

10:07 SX M3 Tip = +150 (absolute Tip=+295) IE=-32 CA=-43

10:07 pupil_collect (round 2)

10:18 SX M3 Tip = -150 (absolute Tip=-5) IE=-37 CA=-68 (but John let some large guider corrections sneak thru)

10:19 pupil_collect (round 3)

10:29 ACQUIRE SX IE=-56 CA=-63 (My pngui doesn't update when Geno makes a change, which solves a mystery between Steve/John.)

10:33 ACTIVE preset again

SX M3 Tip=-300 (absolute Tip =-155) IE=-61 CA=-73

10:xx pupil_collect (round 4)

10:51 SX M3 Tip=-125 (absolute Tip=+20) Tilt=-150 (absolute Tilt=-160) IE=-46 CA=-66

10:5x pupil_collect (round 5)

(adopted!) JMH saved absolute Tip=+20 Tilt=-160 arcsec for LFBG in positions.conf on Tuesday 7-Feb. Pointing model zeropoints for LFBG have been updated accordingly in SX_Gregorian.20131218d.ptmod

After the fact analysis (20201005)
filename              Xavg     Yavg  Radius      pts
pupil_10:51:43.wobble  384.585  214.162  1.88435 4
pupil_10:37:38.wobble  385.610  214.757  5.90179 4
pupil_10:19:15.wobble  385.022  214.955  2.56179 4
pupil_10:07:27.wobble  385.630  215.707  7.14204 4
pupil_09:53:54.wobble  384.587  214.170  4.88169 4
Screenshot at 2020-10-05 20-28-00.png

LFBG Transform Data (attempted)

11:01 ACTIVE preset to StoneJ (RA=13:45) but didn't find star (clouds or pointing?)

11:05 ACQUIRE preset SX IE=-33 CA=-78 DX IE=+87 CA=-213

11:10 ACTIVE preset to StoneJ

11:12 transform_collect - violated copointing limit, so switch to pseudomono without actually figuring out the cause of the violation (maybe a different star in StoneJ).

transform_collect also took an IRTC2 image rather than LUCI

11:14 transform_collect - StoneJ

JMH changed this in transform_collect.pro to get it to park the AGW guide probe for LUCI rather than MODS.
;        gcs_set_gs_sfp, 0.0, -170.0, retval, TEST=test, DEBUG=debug, SILENT=silent
        gcs_set_gs_sfp, 0.0, 150.0, retval, TEST=test, DEBUG=debug, SILENT=silent

11:24 transform_collect, /PAUSE (/PAUSE for taking the LUCI image manually)

luci 0005 good
luci 0006 bad
luci 0007 clouds

transform_collect failed for missing wfsc info.

11:xx transform_collect, /PAUSE StoneJ_small (different star list with different central TS star)

luci 0008 first with lots of clouds
failed after image luci 0010
      IRCL: preset, Stone_J, MODE=GUIDE, GS=1, TS=14, PA=90.0000     (iif_preset)
  
% Compiled module: GCS_STOP_ZERNIKES_TO_PSF.
   => 
% Compiled module: TELESCOPE_STATUS.
% Compiled module: TELESCOPE_STATUS_STRUCT__DEFINE.
% Compiled module: INSTRUMENT_CHECK_ACQUISITION.
% Compiled module: IRC_CHECKSEQIMAGES.
return # = -1
       0
   1

12:00 transform_collect, /PAUSE StoneJ

luci 0011 first image
failed for lack of a star due to clouds

12:09 transform_collect, /PAUSE, start=2 StoneJ
failed because probe can't move to GS=14.

12:12 transform_collect, /PAUSE, start=3 StoneJ
failed because probe can't move to GS=24

12:15 transform_collect, /PAUSE StoneJ_small

luci 0013 first
failed with this message
      IRCL: preset, Stone_J, MODE=GUIDE, GS=1, TS=14, PA=90.0000     (iif_preset)
  
   => 
return # = -1
       0
   1

LUCI1 image parameters: DIT=2.502 NDIT=10 NEXP=1 INTEGRATED BR-gamma LOOKUP=on

12:28 transform_collect, /PAUSE Stone_K3 failed to find star (cloud too thick)

12:31 ACQUIRE preset

12:33 transform_collect, /PAUSE Stone_K2

luci 0016 first image
failed with this message
      IRCL: preset, Stone_K2, MODE=GUIDE, GS=23, TS=19, PA=-45.0000     (iif_preset)
  
   => 
return # = -1
       0
   1

Spent time here waiting for clouds, trying to understand IDL error, and preparing the guidestar catalog for an asteroid. We didn't call Doug again as clouds were already too heavy for successful collection of transform data.

Non-Sidereal Guiding Test (attempted)

nscatalog.py puts catalogs in the wrong place, Needed to do the following cp manually.

cp -p /lbt/catalogs/Scratch/catalogs/coelestina_20170205_1330.catalog /lbt/data/catalogs/lbto/Scratch/catalogs/coelestina_20170205_1330.cat

JMH updated NonSiderealSupport/nscatalog.py to use this path on 20170207.

13:33 Tried a hijacked ACTIVE preset to coelestina, but too bright and too cloudy so we didn't acquire the guide star.

Close

13:35 close for impending dawn

-- %USERSIG{JohnHill - 2017-02-03}%
Topic revision: r11 - 05 Oct 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback